### v1

# Learning Topology-Driven Multi-Subspace Fusion for Grassmannian Deep Networks

Abstract
Grassmannian manifold offers a powerful carrier for geometric representation learning by modelling high-dimensional data as low-dimensional subspaces. However, existing approaches predominantly rely on static single-subspace representations, neglecting the dynamic interplay between multiple subspaces critical for capturing complex geometric structures. To address this limitation, we propose a topology-driven multi-subspace fusion framework that enables adaptive subspace collaboration on the Grassmannian. Our solution introduces two key innovations: (1) Inspired by the Kolmogorov-Arnold representation theorem, an adaptive multi-subspace modelling mechanism is proposed that dynamically selects and weights task-relevant subspaces via topological convergence analysis, and (2) a multi-subspace interaction block that fuses heterogeneous geometric representations through Fréchet mean optimisation on the manifold. Theoretically, we establish the convergence guarantees of adaptive subspaces under a projection metric topology, ensuring stable gradient-based optimisation. Practically, we integrate Riemannian batch normalisation and mutual information regularisation to enhance discriminability and robustness. Extensive experiments on 3D action recognition (HDM05, FPHA), EEG classification (MAMEM-SSVEPII), and graph tasks demonstrate state-of-the-art performance. Our work not only advances geometric deep learning but also successfully adapts the proven multi-channel interaction philosophy of Euclidean networks to non-Euclidean domains, achieving superior discriminability and interpretability.
Introduction
In recent years, unitary subspace modelling on the Grassmannian has achieved great success in tasks which require extended discriminative capacity, such as activity recognition (cherian2017generalized), emotion recognition (liu2014combining), face verification (huang2015projection), and classification of time-series data in brain-computer interfaces (gao2022domain). In the above tasks, Grassmannian typically model the input data with a single and static orthogonal subspace (huang2018building; wang2024grassmannian). However, A fixed single-subspace assumption fails to capture local geometric variations and multi-modal distribution, limiting the expressiveness and adaptability of the model in structurally diverse or heterogeneous task scenarios (hamm2008grassmann). To overcome this issue, we introduce an adaptive multi-subspace representation that dynamically adjusts multiple subspaces to accurately capture task-specific characteristics and the harmonious distribution among subspaces. The convergence of obtaining adaptive subspaces is analysed from a topological perspective (munkres2000topology), satisfying specific demands across different tasks.
In deep neural networks, different vectors in the Euclidean space reflect their specific feature identities. However, the existing Euclidean metric (olver2006applied) driven by the inner product, measures the semantic difference between two vectors, resulting in a distorted relevance in task-specific decisions (as shown in Figure 1 (a) and (d)). Based on non-Euclidean geometry, this distorted relevance can be effectively addressed on Riemannian manifolds. Intuitively, Riemannian manifolds provide compact spaces embedded in the dense Euclidean space, where the accessible path can preserve the geometric consistency (boumal2023introduction). As a typical Riemannian manifold, the Grassmannian (bendokat2024grassmann) can model high-dimensional data as linear subspaces to effectively address semantic distortion. Despite this merit, existing solutions (huang2018building; wang2024grassmannian) are primarily driven by a single subspace, lacking consideration of the collaborative interaction of multi-subspaces, which hinders the characterisation of subspace geometry diversity in terms of extended semantics (as shown in Figure 1 (b) and (e)).
Figure 1: Illustration of the problem evolution. (a) shows semantic distortion of features in Euclidean space. (b) reveals semantic constraint using a single Grassmannian subspace. (c) presents the proposed framework incorporating multi-Grassmannian subspace fusion.
In the Euclidean space, the promising progress of deep learning in the recent decade is attributed to multi-channel interactions and non-linear activations (krizhevsky2012imagenet). In particular, the seminal work LeNet-5 (lecun2002gradient) extracts multi-channel feature maps using stacked convolution blocks, thus obtaining extended representative capacity via channel interaction. Admitting the natural representation power of linear subspace, the potential of multi-subspace interactions has been overlooked in the community. To reveal the power of multi-subspace interaction for the Grassmannian, there are two major issues, i.e., how to perform subspace interaction and how to formulate its deep architecture.
For the first issue, we design a multi-subspace interaction block in a geometric manner, i.e., several intermediate representations are generated by a unified mapping and then fused through alignment, facilitating effective interaction among different Grassmannian subspaces. For the second issue, we leverage an intrinsic geometric approach to ensure that the fusion process occurs at accurate positions on the manifold (tao2024alignmif), which enables the design of our multi-subspace interaction block in a stackable form, allowing for flexible capacity expansion. To the best of our knowledge, this is the first work that introduces deep interactions of Grassmannian subspaces in Riemannian neural networks, as shown in Figure 1 (c) and (f), these deep interactions capture the dynamic relationships between multiple subspaces, effectively modelling their geometric structures, semantic features, and mutual influences. Euclidean deep networks naturally implement backpropagation, given that the gradient superposition property satisfies the axioms of linear spaces (fei2025survey). However, parameter optimisation on a Riemannian manifold must be strictly restricted to its tangent space  (smith2014optimization). To this end, we develop a topology-driven framework for adaptive multi-subspace construction on the Grassmannian. In summary, our main contributions are the following:
  • We propose an adaptive multi-subspace modelling approach tailored to diverse recognition tasks.
  • We propose a novel Grassmannian Multi-Subspace Fusion network (GMSF-Net) to fuse heterogeneous subspace representations.
  • We propose a topology-driven framework with theoretical guarantees for adaptive Grassmannian subspaces construction.
Related Work
The Grassmannian is the set of all linear subspaces of a fixed dimension, showing notable algebraic and geometric merits, reflecting the stability and invariance of subspace geometric structures (souza2020interface). In related research, deep learning applications on the Grassmannian have evolved from shallow methods to deep models.
In the early research stage, shallow Grassmannian learning (zhang2018grassmannian) is achieved by projecting subspaces onto the manifold, such as discriminant analysis (baudat2000generalized), high-dimensional data clustering (von2007tutorial), and low-rank matrix completion (dai2012geometric). However, shallow learning on the manifold cannot mine deep geometric characteristics from the data, thus only suitable for simple tasks. The development of GrNet breaks this limitation by formulating a deep network version of the Grassmannian (huang2018building). In particular, GrNet projects data onto the Grassmannian and follows the manifold’s geometric properties to design network layers such as FRMap, OrthMap, and ProjMap. A brief summary of GrNet is provided in Appendix B. Typically, GrNet considers the manifold’s constraints and the low-dimensional subspace representation of the data, thereby enhancing the model capacity to learn delicate geometric structures and effectively solve nonlinear relationships and high-dimensional problems.
In recent studies, equipped with the self-attention mechanism, models can focus on the global relevance of the input and build long-range dependencies among them, greatly enhancing the model capacity (vaswani2017attention). To this end, to explore the potential of self-attention on the Grassmannian, GDLNet (wang2024grassmannian) integrates the mechanism to build dependencies between different subspaces. It helps the model effectively focus more on important subspace information. However, GrNet and GDLNet typically adopt a single and static subspace to model the original global features, which makes it difficult to extract diverse geometric subspaces, thus delivering internal disturbances and weaker discriminative power. To address this issue, we introduce an adaptive multi-subspace construction mechanism to capture coherent geometric features. Additionally, we design a multi-subspace interaction block to fuse different subspaces, facilitating intra-subspace transformation and inter-subspace complement.
Preliminaries
This section provides a brief overview of the basic geometries of Grassmannian and Topology. More detailed review and relevant notations can be found in Appendix A.
Grassmannian
Grassmannian geometries: The Grassmannian $\mathcal{G}(n,p)$ is the set of all $p$-dimensional linear subspaces of $\mathbb{R}^{n}$. Each point represents a $p$-dimensional subspace, typically represented by an orthonormal basis matrix $U$. The tangent space, exponential map, and logarithmic map are key tools to understand its geometric structure, and these concepts will serve as prerequisites for subsequent operations (a detailed explanation is provided in Appendix A.2).
Grassmannian Metric: Points on the Grassmannian represent different subspaces, and their distances are typically measured by metrics such as the Projection metric (huang2015projection) or Principal Angles metric (qiu2005unitarily). Considering computational efficiency, this paper adopts the Projection metrics to analyse points on the manifold. For two subspaces $X_{1}$ and $X_{2}$, the Projection metric is defined as:
$\mathit{d_{p}}(\mathit{X}_{1},\mathit{X}_{2})=2^{-1/2}\|\mathit{X}_{1}\mathit{X}_{1}^{T}-\mathit{X}_{2}\mathit{X}_{2}^{T}\|_{F},$ (1)
where $\|\cdot\|_{F}$ is the Frobenius norm (harandi2013dictionary). This metric serves as an effective tool for quantifying the distance between subspaces and enabling geometric analysis on the Grassmannian (fei2025survey).
Topology
Topological spaces underpin Riemannian geometry by abstracting spatial properties beyond specific shapes (lee2000introduction). On the Grassmannian, each subspace inherently constitutes a topological space. This topological framework rigorously characterises subspace structure and metric, ensuring the critical convergence property for sequences of subspaces. This theoretical guarantee of convergence motivates the design of adaptive multi-subspace modelling, ensuring stable convergence towards limiting subspaces during optimisation and enabling efficient geometric feature extraction and fusion. This section provides a brief introduction to the topological structure, with detailed definitions and proofs presented in Appendices A.3 and F.
Figure 2: Overview of GMSF-Net. (a) Architecture of GMSF-Net. (b) Structure of the Adaptive Multi-Subspace Encoder (AMSE). (c) Fine-Grained Design. The proposed efficient Adaptive Multi-Subspace Construction (AdaMSC). (d) Subspace Interaction Design. The proposed discriminative Multi-Subspace Interaction Block.
Approach
In this section, we will systematically introduce our method. We first introduce the adaptive multi-subspace encoder on the Grassmannian, analysing its design principles in depth and proving its convergence based on topological theory. Subsequently, we explain the design of the multi-subspace interaction block based on a geometric approach in detail. Figure 2 (a) presents an overview of GMSF-Net.
Adaptive Multi-Subspace Encoder (AMSE)
This subsection introduces the Adaptive Multi-Subspace Encoder (AMSE) on the Grassmannian, which adaptively constructs multiple coherent subspaces to preserve the diverse geometric characteristics of data, as illustrated in Figure 2 (b). This subsection introduces the construction of adaptive multiple subspaces and the analysis of their convergence in detail.
Adaptive Multi-Subspace Construction (AdaMSC)
The construction of adaptive multi-subspaces is the core of AMSE, as shown in Figure 2 (c), which builds multiple stable and harmonious low-dimensional subspace representations on the high-dimensional Grassmannian space $\mathcal{G}(n,k)$, and adaptively adjusts these subspaces based on different tasks. In particular, we first extract frame-level features and compute a covariance matrix $X\in\mathbb{R}^{n\times n}$ to model the statistical dependencies of features along the temporal dimension (nguyen2019neural). Subsequently, we apply the Schmidt orthogonalisation process to map $X$ into a set of low-dimensional orthogonal subspaces (yaghooti2024gram):
$\mathcal{S}=\{S_{1},S_{2},\dots,S_{k}\},\quad S_{i}^{\top}S_{j}=0\quad(\forall\,i\neq j),$ (2)
where each subspace $S_{j}\in\mathcal{G}(n,1)$. This design aims to decouple statistical information from physical information in the features, transforming them into a set of localised geometric encodings that facilitate subsequent subspace modeling (edelman1998geometry).
We define the index set of subspaces as $\mathcal{I}=\{i_{1},\dots,i_{k}\}$, where each index $i_{j}$ corresponds to a subspace $S_{j}\in\mathcal{G}(n,1)$. For each new subspace $S^{\prime}_{m^{\prime}}$, we initialise a set of learnable weight parameters $\mathcal{W}^{(m^{\prime})}=\{w_{1}^{(m^{\prime})},\dots,w_{k}^{(m^{\prime})}\}$, which are normalised via the Softmax function to obtain the weight distribution $\widetilde{\mathcal{W}}^{(m^{\prime})}=\{\widetilde{w}_{1}^{(m^{\prime})},\ldots,\widetilde{w}_{k}^{(m^{\prime})}\}$. Each $\widetilde{w}_{j}^{(m^{\prime})}$ reflects the relative importance of the atomic subspace $S_{j}$ during the construction of $S^{\prime}_{m}$. We then select the top $p$ indices with the highest values in $\widetilde{\mathcal{W}}^{(m^{\prime})}$, obtaining the index set $\{i_{j_{1}},i_{j_{2}},\ldots,i_{j_{p}}\}$ and the corresponding weights $\{\widetilde{w}_{j_{1}}^{(m^{\prime})},\ldots,\widetilde{w}_{j_{p}}^{(m^{\prime})}\}$. The $m^{\prime}$-th new subspace is constructed from these weighted subspaces, with the construction defined as above based on Theorem F.1 in the Appendix:
$S^{\prime}_{m^{\prime}}=[\widetilde{w}_{j_{1}}^{(m^{\prime})}S_{j_{1}},\ \widetilde{w}_{j_{2}}^{(m^{\prime})}S_{j_{2}},\ \dots,\ \widetilde{w}_{j_{p}}^{(m^{\prime})}S_{j_{p}}].$ (3)
In this way, the new subspace $S^{\prime}_{m^{\prime}}\in\mathcal{G}(n,p)$ is composed of distinct key atomics which enable efficient modelling and adaptive adjustment of diverse features and lay a solid foundation for subsequent multi-subspace fusion. The detailed process is shown in algorithm 1.
Algorithm 1 Adaptive Multi-Subspace Construction
1:Input: Low-dimensional subspaces $\{S_{i}\}_{i=1}^{k}$, $S_{i}\in\mathcal{G}(n,1)$, subspace dimension $p$, number of output subspaces $m^{\prime}$
2:Parameter: Importance matrix $\mathcal{W}\in\mathbb{R}^{m^{\prime}\times k}$
3:Output: Weighted multi-subspaces $\{S^{\prime}_{i}\}_{i=1}^{m^{\prime}}$
[…]
5:for $i=1$ to $m^{\prime}$ do
6:  Sort indices: $I_{i}\leftarrow argsort_{descending}(\mathcal{W^{\prime}}[i,:])$
7:  Select top-$p$ components: $\widetilde{S}_{i}\leftarrow S_{i}[:,I_{i}[1:p]]$
8:  Get corresponding weights: $\mathbf{w}_{i}\leftarrow\mathcal{W^{\prime}}[i,I_{i}[1:p]]$
9:  $T_{i}\leftarrow\widetilde{S}_{i}\odot(\mathbf{1}_{n}\cdot\mathbf{w}_{i}^{\top})$ $\mathbf{1}_{n}$: all-ones vector}
10:end for
11: Concatenate subspaces: $\{S^{\prime}_{i}\}_{i=1}^{m^{\prime}}\leftarrow\bigcup_{k=1}^{m^{\prime}}\{T_{k}\}$
Next, we introduce the mutual information function $f(S^{\prime})$ to measure the complementarity between different new subspaces $S_{i}^{\prime}$ and $S_{j}^{\prime}$ (liu2009feature). This function aims to enhance the discriminative ability of the overall representation by maximising the information complementarity between subspaces:
$f(S^{\prime})=I(S_{i}^{\prime},S_{j}^{\prime}),\quad i\neq j,\quad i,j\in\{1,\dots,m^{\prime}\}.$ (4)
The specific formula for the mutual information $I(S_{i}^{\prime},S_{j}^{\prime})$ is:
$I(S_{i}^{\prime},S_{j}^{\prime})=\int\int p(S_{i}^{\prime},S_{j}^{\prime})\log\frac{p(S_{i}^{\prime},S_{j}^{\prime})}{p(S_{i}^{\prime})p(S_{j}^{\prime})}\,dS_{i}^{\prime}\,dS_{j}^{\prime}.$ (5)
In practical modelling, since it is difficult to directly obtain the probability distributions $p(S_{i}^{\prime})$ and $p(S_{j}^{\prime})$ for the subspaces, we typically use embedding or feature-based estimation methods to approximate the mutual information (for detailed information see the Optimisation Strategies(maeda2025estimation).
Analysis of Adaptive Subspace Convergence
We further analyse multi-subspace construction on the Grassmannian from a topological perspective to ensure model convergence. By introducing a metric topology and projection metric, we ensure the subspaces converge iteratively. In the metric topology, the distance on the Grassmannian is defined by the projection metric, as shown in Equation (1). This metric satisfies non-negativity, symmetry, and the triangle inequality, ensuring that the Grassmannian is a metric topology. The proof is in Appendix F.2.
In this metric topology, the convergence of a sequence of subspaces is naturally defined as detailed in Proposition F.3 in the Appendix. During training (ionescu2015training), the subspace is iteratively updated through Riemannian gradient descent, with the update rule as follows:
$S^{\prime}(t+1)=\text{Exp}_{S^{\prime}(t)}\left(-\eta\,\text{Proj}_{T_{S^{\prime}(t)}}\nabla L(S^{\prime})\right),$ (6)
where $\text{Exp}_{S^{\prime}(t)}$ is the exponential map, $\text{Proj}_{T_{S^{\prime}(t)}}$ is the tangent space projection, $\nabla L(S^{\prime})$ is the gradient,and the step size $\eta$ controls the learning rate. The proof is in Appendices F.3 to F.5. Due to the continuity and differentiability of the loss function with a lower bound, along with the completeness of the manifold and the continuity of the exponential map (eschenauer2001topology), this subspace eventually satisfies (see Appendix F.6):
$d(S^{\prime}(t),S^{*})\to 0\quad\text{as}\quad t\to\infty.$ (7)
This means that in the topology induced by the projection metric, the sequence converges to the stable subspace $S^{*}$, ensuring the convergence and consistency of the model’s geometric modelling process.
Multi-Subspace Interaction Block
The multi-subspace interaction block fuses multiple Grassmannian subspaces into a single integrated subspace, as shown in Figure 2, effectively capturing geometric features of data and enhancing the ability to represent complex geometric cues. This block consists of two main components: Grassmannian Multi-Representations and Subspace Interaction.
This work introduces a semantic perspective for multi-subspace interaction design, addressing the often-overlooked intuition behind such interactions. We posit that distinct subspaces, governed by different geometric frameworks, capture complementary task-relevant information through diverse local geometric features. Their interaction transcends mere feature fusion, embodying semantic co-aggregation within the manifold. By incorporating alignment constraints and weighted aggregation, our model dynamically integrates multi-subspace representations, enhancing both discriminability and geometric diversity.
Grassmannian Multi-Subspace Representations (GMSR)
Grassmannian Multi-Subspace Representations (GMSR) generates subspace representations with distinct geometric features under a unified geometric framework by applying the same learnable matrix to a set of input subspaces $\{S^{\prime}_{m^{\prime}}\}$, producing a set of output representations $\{X_{GMSR}^{c,m^{\prime}}\}$. The formulation is as follows:
$X_{GMSR}^{c,m^{\prime}}=f_{GMSR}^{(c)}(S^{\prime}_{m^{\prime}},W_{c})=W_{c}^{T}S^{\prime}_{m^{\prime}},$ (8)
where $S^{\prime}_{m^{\prime}}\in\mathcal{G}(n,p)$ is the $m$-th output subspace from the AMSE module, $W_{c}\in\mathbb{R}^{n\times m},(m<n)$ is the mapping matrix that is basically required to be a full rank matrix (huang2017riemannian), $X_{GMSR}^{c,m^{\prime}}\in\mathcal{G}(m,p)$ is the new Grassmannian representations. Generally, $W_{c}^{T}S^{\prime}$ is not an orthonormal basis matrix, we employ QR or SVD decomposition to preserve the intrinsic geometric structure of the data. To obtain a closed-form expression for the mapping matrix $W_{c}$ during the optimisation process, we constrain it to the Stiefel manifold $St(n,m)$ (huang2017riemannian). We define $c$ geometric frameworks as mapping matrices $\left\{W_{1},\ldots,W_{c}\right\}$ capturing different subspace representations, where $c$ denotes the number of geometric frameworks. The learnability of the mapping matrix enables the model to adaptively adjust the geometric framework, capture the geometric information of multiple subspaces, and improve modelling capacity for geometric feature extraction and subspace fusion.
Grassmannian Subspace Interaction (GSI)
In the GMSR layer, we construct subspace representations under various geometric frameworks $c$. The input is a set of $m^{\prime}$ Grassmannian subspace points denoted as $\{P_{i}\}_{i=1}^{m^{\prime}}$, where each $P_{i}=X_{GMSR}^{c,i}\in\mathcal{G}(m,p)$. To enable effective fusion of the set $\{P_{i}\}_{i=1}^{m^{\prime}}$, we introduce the Fréchet mean to preserve the intrinsic geometric structure (pennec2006riemannian). Specifically, Fréchet mean $\Gamma$ is computed as the point that minimises the sum of squared distances on the manifold (yang2010riemannian):
$\Gamma\!=\!\mathcal{F}_{w}(\{P_{i}\}_{\!i\leq m^{\prime}})\!:=\!\textit{arg}\min_{\!\Gamma\in\mathcal{G}(m,p)}\sum_{i=1}^{m^{\prime}}\!w_{i}\delta_{R}^{2}(\Gamma,P_{i}),$ (9)
where, $w_{i}$ denotes the weight of each point, satisfying $w_{i}\geq 0$ and $\sum_{i=1}^{m^{\prime}}w_{i}=1$. The function $\delta_{R}(\cdot,\cdot)$ represents the geodesic distance on the Grassmannian.
When $m^{\prime}=2$, i.e., the weights are given by $\{w,1-w\}$, a closed-form solution exists, which exactly corresponds to the geodesic between the two points $P_{1}$ and $P_{2}$, parameterised by $w\in[0,1]$:
$\Gamma=V\left[P_{1}\cos(w\Theta)+U\sin(w\Theta)\right]V^{\top},$ (10)
where $P_{1}^{\top}P_{2}=V\cos(\Theta)Q^{\top}$ is part of its singular value decomposition (SVD), $U,\Theta,V$ come from the SVD of $(I-P_{1}P_{2}^{\top})P_{2}V$, and the parameter $w\in[0,1]$ controls the position of the fused point (beik2021learning).
When $m^{\prime}>2$, the optimisation problem does not have a closed-form solution. Therefore, the Karcher flow algorithm is employed to perform optimisation in the tangent space iteratively, ultimately converging to the Fréchet mean through exponential mapping. The update formula of the Karcher flow (karcher1977riemannian; yang2010riemannian) algorithm is as follows (Appendix C provides detailed steps):
$\Gamma_{new}=\exp_{\Gamma}\left(\Gamma,\alpha\cdot\frac{1}{m^{\prime}}\sum_{i=1}^{m^{\prime}}\log_{\Gamma}(\Gamma,P_{i})\right),$ (11)
where $\alpha$ is the step size, $\Gamma$ is the current mean, and $P_{i}$ represents each subspace within the same geometric framework. $\log_{G}(\cdot,\cdot)$ and $\exp_{G}(\cdot,\cdot)$ are provided in the Appendix A.
Deep Subspace Interaction and Optimisation Principles
Deep Subspace Interaction Principles: We propose a stackable GMSF-Net by applying multiple blocks to meet the flexible and diverse requirements of practical applications. The core principle is to constrain the feature manifold distribution through Riemannian batch normalisation, thereby enhancing geometric discriminability while stabilising the feature space topology. The feature construction based on Riemannian BN provides navigation signals with geometric interpretability, enabling the stacking structure to systematically enhance and refine the features.
For the input of the AMSE block, i.e. statistical features, we introduce a Riemannian normalisation strategy (brooks2019riemannian) based on manifold geometry, mapping statistical features to the SPD manifold (nguyen2019neural) and normalising them to extract more discriminative key Riemannian features (see Appendix E for the detailed implementation of SPDBN):
$\hat{X}=\exp_{\hat{B}}\left(\hat{B}^{1/2}\left((\hat{B}\!B)^{-1/2}X(\hat{B}\!B)^{-1/2}\right)\hat{B}^{1/2}\right),$ (12)
where $\hat{X}$ is the refined feature, $X$ is the original input, $B$ is the Riemannian barycenter, and $\hat{B}$ is the learnable common feature, which is iteratively updated using the Riemannian geodesic formula (yair2019parallel).
Optimisation Strategies: In order to effectively train the proposed GMSF-Net, we restructure the loss function by combining classification cross-entropy loss (hinton1995wake) and mutual information loss, effectively capturing diverse geometric subspaces. The total loss function for the proposed method is as follows:
$\mathcal{L}_{total}=\mathcal{L}_{CE}+\lambda\cdot\mathcal{L}_{C},$ (13)
where, $\mathcal{L}_{total}$ is the total loss, $\mathcal{L}_{CE}$ is the cross-entropy loss, $\mathcal{L}_{C}$ is the mutual information loss, and $\lambda$ is a hyperparameter that balances classification accuracy with structural preservation. The details of these two loss functions (Appendix D.1) and their backpropagation derivations (Appendix D.2) are provided in Appendix D.
Evaluation
In this section, we evaluate the performance of the proposed GMSF-Net on two challenging classification tasks: HDM05 (muller2007mocap) and FPHA (garcia2018first) datasets for video-based 3D action recognition, and the MAMEM-SSVEP-II (pan2022matt) dataset for EEG signal classification. Additionally, we perform node classification (NC) and link prediction (LP) tasks on graph datasets, where the underlying structure of the data can naturally be represented as Grassmannian, possessing multi-subspace geometric properties. More experimental details are presented in Appendix G.
Method Acc.(%) Size(MBMB) FLOPs($M$)
SPDNet 87.65%±1.02 13.60 1595.50
GrNet 78.79%±1.82 6.73 38.60
MATT 87.70%±0.68 1.83 142.07
SPDNetBN 89.33%±0.49 13.63 1902.97
GDLNet 87.60%±0.69 1.83 33.69
GMSF-Net-1Block 90.43%±0.74 1.20 48.42
GMSF-Net-2Blocks 90.70%±0.70 1.26 63.44
GMSF-Net-3Blocks 91.22%±0.53 1.30 81.07
Table 1: Comparison Between GMSF-Net and Other Riemannian Solutions on the FPHA Dataset.
Method Acc.(%) Size(MBMB) FLOPs($M$)
SPDNet 60.45%±1.12 11.71 2050.49
GrNet 59.23%±1.78 6.88 81.96
MATT 62.25%±1.68 3.99 442.45
GDLNet 60.08%±1.78 3.95 89.11
GMSF-Net-1Block 63.64%±1.24 3.33 96.08
GMSF-Net-2Blocks 63.98%±1.07 3.34 100.07
GMSF-Net-3Blocks 64.19%±0.88 3.39 114.12
Table 2: Comparison Between GMSF-Net and Other Riemannian Solutions on the HDM05 Dataset.
Method Acc.(%) Size(MBMB) FLOPs ($M$)
EEGNet 53.72%±7.23 0.075 60.13
ShallowConvNet 56.93%±6.97 0.18 127.00
EEG-TCNet 55.45%±7.66 0.016 60.80
SCCNet 62.11%±7.70 0.55 108.56
MBEEGSE 56.45%±7.27 1.64 52.40
FBCNet 53.09%±5.67 0.24 51.80
SPDNet 62.30%±3.12 2.81 2274.56
GrNet 61.23%±3.56 1.95 2020.78
MATT 65.19%±3.14 1.97 2068.44
SPDNetBN 62.76%±3.01 2.81 2312.78
GDLNet 65.52%±2.86 1.95 2028.34
GMSF-Net-1Block 66.74%±1.79 1.94 2020.14
GMSF-Net-2Blocks 66.32%±1.84 1.94 2021.78
GMSF-Net-3Blocks 66.87%±1.46 1.94 2032.55
Table 3: Performance comparison between GMSF-Net and the baseline on SSVEP.
Dataset Disease Airport PubMed CoRA
Task LP NC LP NC LP NC LP NC
Euc 59.8%±2.0 32.5%±1.1 92.0%±0.0 60.9%±3.4 83.3%±0.1 48.2%±0.7 82.5%±0.3 23.8%±0.7
Hyp 63.5%±0.6 45.5%±3.3 94.5%±0.0 70.2%±0.1 87.5%±0.1 68.5%±0.3 87.6%±0.2 22.0%±1.5
Euc-Mixed 49.6%±1.1 35.2%±3.4 91.5%±0.1 68.3%±2.3 86.0%±1.3 63.0%±0.3 84.4%±0.2 46.1%±0.4
Hyp-Mixed 55.1%±1.3 56.9%±1.5 93.3%±0.0 69.6%±0.1 83.8%±0.3 73.9%±0.2 85.6%±0.5 45.9%±0.3
MLP 72.6%±0.6 28.8%±2.5 89.8%±0.5 68.6%±0.6 84.1%±0.9 72.4%±0.2 83.1%±0.5 51.5%±1.0
HNN 75.1%±0.3 41.0%±1.8 90.8%±0.2 80.5%±0.5 94.9%±0.1 69.8%±0.4 89.0%±0.1 54.6%±0.4
GMSF-Net-1Block 95.0%±1.0 79.7%±2.9 93.4%±0.2 82.5%±0.8 94.6%±0.1 72.7%±0.5 88.6%±0.6 57.8%±1.2
GMSF-Net-2Blocks 94.6%±0.7 81.1%±1.9 92.7%±0.6 80.7%±1.0 95.0%±0.1 72.8%±0.4 89.3%±0.6 55.8%±1.5
GMSF-Net-3Blocks 95.5%±0.4 82.7%±0.8 91.2%±0.2 78.3%±0.6 93.6%±0.3 72.9%±0.3 88.9%±0.3 56.2%±1.1
Table 4: Performance comparison between GMSF-Net and the baselines on Graph Tasks.
With Interaction Without Interaction
Configuration HDM05 FPHA SSVEP HDM05 FPHA SSVEP
Adaptive-Subspace 63.64%±1.24 90.43%±0.74 66.74%±1.79 56.49%±1.37 80.68%±2.15 59.83%±2.72
Random-Subspace 50.29%±2.10 72.47%±3.19 56.01%±2.46 52.50%±1.82 73.10%±2.19 57.43%±1.10
Fixed-Subspace 53.04%±0.88 83.06%±0.61 66.05%±1.95 51.12%±3.55 76.67%±0.16 58.85%±1.19
Table 5: Performance Comparison of Subspace Construction and Interaction Mechanisms.
Experimental Evaluation on the Classification Datasets
We evaluate the proposed GMSF-Net on video-based 3D action recognition and EEG signal classification tasks. For action recognition, we use the FPHA and HDM05 datasets, following the experimental settings of SymNet (wang2021symnet) and GrNet (huang2018building). For EEG classification, we follow the preprocessing pipeline of GDLNet (wang2024grassmannian) and conduct experiments on the MAMEM-SSVEP-II dataset. In the model evaluation, we compare performance involving different numbers of multi-subspace interaction blocks. As performance saturates with more blocks, we report results with up to three blocks in this section. For each task, we highlight the best performance results in bold.
FPHA: Following (wang2021symnet), we evaluate GMSF-Net using three network configurations: 1Block, 2Blocks, and 3Blocks. The results from 10 random experiments (mean ± std) are shown in Table 1. Notably, GMSF-Net consistently outperforms most Riemannian deep models, with a 12.43% improvement over GrNet. As the number of blocks increases, performance improvements plateau, yet the effectiveness of the adaptive subspace interaction method remains evident. Additionally, our model demonstrates significant advantages in terms of model size and computational cost.
HDM05: Following (huang2018building), we also adopt three architectures to evaluate GMSF-Net: 1Block, 2Blocks, and 3Blocks configurations. The 10-fold results (mean ± std) are presented in Table 2. It should be noted that GMSF-Net achieves a performance improvement of up to 4.96% over GrNet. This highlights the effectiveness of our approach and significantly reduces memory and computational costs compared to SPD manifold-based neural networks. Results on HDM05 and FPHA demonstrate that GMSF-Net consistently outperforms other methods in 3D action recognition.
MAMEM-SSVEP-II: In Table 3, we follow the protocol of GDLNet and report the results of ten random cross-validation runs (mean ± std). It can be observed that GMSF-Net outperforms existing Riemannian models as well as EEG deep learning models, including EEGNet (lawhern2018eegnet), ShallowConvNet (schirrmeister2017deep), EEG-TCNet (ingolfsson2020eeg), SCCNet (wei2019spatial), MBEEGSE (altuwaijri2022multi), and FBCNet (mane2021fbcnet), in terms of overall performance. The accuracy is improved by 5.64% over GrNet, and by 1.68% and 1.35% over the Riemannian self-attention models MATT (pan2022matt) and GDLNet, with model stability also enhanced. EEG visualisation results are shown in Appendix G.4. These results further highlight the versatility and effectiveness of our framework.
Experimental Evaluation on Graph Datasets
To evaluate the broad applicability of GMSF-Net, we perform node classification (NC) and link prediction (LP) on graph datasets with prominent Grassmannian subspace structures. The implementation details is provided in Appendix G.2.2. The experimental results show that increased deviation from the ideal Grassmannian structure leads to performance degradation due to modelling limits on non-natural manifold geometries. In contrast, GMSF-Net exhibits greater performance gains on datasets more aligned with the Grassmannian assumption, validating the soundness and adaptability of its design. Specifically, GMSF-Net achieves notable gains on Cora (sen2008collective) and Disease (chami2019hyperbolic) datasets, which exhibit clear subspace structures, while improvements are more modest on structurally complex PubMed (sen2008collective) and Airport (chami2019hyperbolic) datasets that deviate from ideal subspace assumptions. GMSF-Net achieves notable gains on all graph datasets except PubMed, as shown in Table 4.
Ablation Experiments and Analysis
As shown in Table 5, we conducted six ablation studies across three datasets to evaluate the effectiveness of the adaptive subspace and interaction mechanisms. The results show that the framework combining adaptive subspace construction with the interaction mechanism consistently achieves the best performance (HDM05: 63.64%, FPHA: 90.43%, SSVEP: 66.74%), significantly outperforming other configurations. This validates the necessity of adaptive mechanisms for optimising the feature space and highlights the synergistic enhancement provided by the interaction mechanism within high-confidence subspaces. As shown in Table 5, the random subspace achieves the lowest performance (e.g., HDM05: 52.50% $\rightarrow$ 50.29%), indicating that low-quality interactions may introduce noise. While the fixed subspace reaches 83.06% on FPHA, its static structure limits its ability to adapt to diverse geometric characteristics. The adaptive subspace is essential for effectively collaborating with the interaction mechanism to capture feature correlations and enhance generalisation performance. Additional experiments on model hyperparameters are provided in Appendix G.3.
Conclusion
This study innovatively proposes a topology-driven multi-subspace fusion framework on the Grassmannian, unifying adaptive geometric representation learning with differentiable subspace interaction. By dynamically optimising subspaces to generate task-oriented geometric representation primitives and incorporating a geometry-aware interaction network for subspace topology correlation modelling, we successfully achieve simultaneous improvements in model accuracy and noise robustness in cross-dataset experiments. This work provides a new differentiable manifold learning paradigm for fields such as computer vision and natural language processing.